OpenStudy (anonymous):

what are the factors of(2x+3)^2 - (4y-5)^2???

OpenStudy (anonymous):

Treat that as a difference of two squares. \[a^2-b^2 = (a+b)(a-b)\] where: \[a = 2x+b, b = 4y-5\]

OpenStudy (anonymous):

This is a difference of two squares that returns the product: a^2 - b^2 = (a-b)(a+b) Let a = 2x + 3 and b = 4y - 5 a-b = 2x + 3 - 4y + 5 = 2x - 4y + 8 a + b = 2x + 3 + 4y - 5 = 2x + 4y - 2 The factors are: (2x - 4y + 8)(2x + 4y - 2)